# Are Dense Labels Always Necessary for 3D Object Detection from Point Cloud?

### v1

@@ 3 Preliminaries
where $E_{i}=\{E_{ij}\},j=1,\cdots,N_{i}$ is a set of detected 3D objects in the $i$-th scene with an input scene $P_{i}\in\mathbb{R}^{N\times 3}$ containing $N$ points, $\phi$ is the weight parameter of $F$, $M$ is the total number of point cloud scenes in the dataset, and $E_{ij}$ can be written as
@@ 3.2 Fully-Supervised 3D Object Detection
where classification loss $\mathcal{L}_{cls}$ is used to learn the classifier to predict the category of a 3D object, and $\mathcal{L}_{reg}$ is utilized to learn the location, size and heading angle of the 3D object. If we take the proposal-based two-stage 3D detector Voxel-RCNN [23] as an example, the Eq. (7) degenerates as
@@
where $\mathcal{L}_{cls}^{RPN}$ is the Focal Loss [7] for classification in the region proposal network (RPN) stage, $\mathcal{L}_{cls}^{Head}$ is the Binary Cross Entropy Loss for confidence prediction in the detection head stage, and Huber Loss is exploited in $\mathcal{L}_{reg}^{RPN}$ and $\mathcal{L}_{reg}^{Head}$ for box regression.
@@ 4 Sparsely-Supervised 3D Object Detection Framework and Learning Algorithm
Though fully-supervised 3D object detection obtains impressive performance, they require large-scale, precisely densely-annotated 3D training data. This data collection process is quite time-consuming and labor-intensive, which prevents the potential applicability of 3D object detection in real-world settings. In this work, we study the following sparsely annotated 3D object detection setting, i.e., we have only access to one 3D bounding box annotation $\tilde{G}_{i}=\{(b_{i1},y_{i1})\}$ for any scene $P_{i}$, and the annotations of remaining object instances are missing. Such inexact and incomplete supervision greatly saves annotation efforts. In contrast to prior weakly-supervised attempts [59, 24, 11] (as shown in Fig. 1), our novel sparse supervision form introduces more potential practical and commercial benefits, while bringing some new challenges (please refer to Section 1). In this section, we will present the proposed SS3D++ method to address these challenges.
@@ 4.1 Overview SS3D++ Framework and Modeling
denote three different curriculum learning objectives,
$\mathcal{L}_{s}=\mathcal{L}_{F}(E_{i1}(\phi),\tilde{G}_{i1}),$ (14)
$\mathcal{L}_{p}=\sum_{j=2}^{N_{i}}\sum_{c=1}^{C}u_{ijc}v_{ijc}k_{ijc}\mathcal{L}^{c}_{F}(E_{ij}(\phi),\tilde{U}_{ij}),$ (15)
@@
$u^{(t+1)}_{ijc}$ =$=\left\{\begin{array}[]{ll}1&\mathcal{L}^{c}_{cls}(p_{ij},y_{ij}^{(t)};\phi^{(t)})<\lambda^{u}_{ic}\\ 0&\mathcal{L}^{c}_{cls}(p_{ij},y_{ij}^{(t)};\phi^{(t)})\geq\lambda^{u}_{ic}\end{array},\right.$ (18)
$v^{(t+1)}_{ijc}$ =$=\left\{\begin{array}[]{ll}1&\mathcal{L}^{c}_{reg}(\mathcal{A}(r_{ij}),\tilde{r}_{ij};\phi^{(t)})<\lambda^{v}_{ic}\\ 0&\mathcal{L}^{c}_{reg}(\mathcal{A}(r_{ij}),\tilde{r}_{ij};\phi^{(t)})\geq\lambda^{v}_{ic}\end{array},\right.$ (21)
$k^{(t+1)}_{ijc}$ =$=\left\{\begin{array}[]{ll}1&d^{c}(r_{ij};\phi^{(t)})<\lambda^{k}_{ic}\\ 0&d^{c}(r_{ij};\phi^{(t)})\geq\lambda^{k}_{ic}\end{array},\right.$ (24)
@@
${\phi}_{s}^{*}=\mathop{\arg\min}_{{\phi}}(\mathcal{L}_{s}+\mathcal{L}_{p}).$ (26)
$\phi_{m}\leftarrow\alpha\phi_{m}+(1-\alpha)\phi_{s},$ (28)
[…]
Algorithm 1 Algorithm of Our SS3D++ Method.
@@ 4.2.2 Adaptive Sample Selection Curriculum Setups
Algorithm 2 Algorithm of Confident Missing-Annotated Instance Mining.
@@
3:     Obtain ${E}_{i}$ and $\hat{{E}}_{i}$ by inputting $P_{i}$ and $\hat{P}_{i}$ to the detector $F_{m}(\cdot;\phi^{(t)})$ based on Eq. 5;
4:     Set $\lambda^{u}_{ic},\lambda^{v}_{ic},\lambda^{v}_{ic}$ by adaptive sample selection curriculum setups in Section 4.2.2;
5:     Perform joint multi-criteria sample selection process via Eq. 18, Eq. 21 and Eq. 24;
6:     Generate high quality of pseudo annotations $\tilde{U}_{i}$ for selected instances via Eq. 25;
@@
Figure 7: Illustration of the proposed reliable background mining module. To start, we feed the original point cloud to the detector without NMS to produce near-duplicate predictions, and leverage the instances stored in the instance bank to filter out unreliable 3D object points. This will lead to broken scene with reliable background, which is further processed through the GT sampling data augmentation in Section 4.4.
@@ 4.3 Reliable Background Mining
As illustrated in Fig. 7, we construct an instance bank which initially just stores the sparsely annotated instances and then progressively stores mined highly confident pseudo instances (please see in Section 4.2). Specifically, we denote the predicted 3D bounding boxes as $R_{i}$, and the bounding boxes of the scene $P_{i}$ in instance bank $\mathcal{B}$ as $\mathcal{B}_{i}$, which contains the sparsely annotated instance and previously mined pseudo instances. And then we delete all points of the predicted bounding boxes $R_{i}$ in the scene $P_{i}$ to get the confident scene $\tilde{P}_{i}$. Next, the bounding boxes and corresponding points stored in the instance bank $\mathcal{B}_{i}$ of the $i$-th scene will fill into the scene $P_{i}$. The instance bank helps to avoid wrongly discarding reliable 3D objects whilst removing the ambiguous points for each training scene.
@@
Algorithm 3 Algorithm of Reliable Background Mining.
@@ 4.4 Confident Fully-Annotated Scene Generation
Method Data 3D Detection (Car) BEV Detection (Car) 3D Detection (Cyc.) BEV Detection (Cyc.) Average
Easy Mod Hard Easy Mod Hard Easy Mod Hard Easy Mod Hard
SS3D (PointRCNN-based [14]) Sparse 87.18 77.10 76.13 89.74 87.41 85.71 86.62 73.22 66.92 87.21 74.27 71.54 80.25
SS3D++ (PointRCNN-based [14]) Sparse 88.63 77.90 77.13 90.26 87.65 86.10 86.73 74.81 70.28 87.76 74.64 71.81 81.14
SS3D (PV-RCNN-based [6]) Sparse 89.49 79.30 78.28 90.45 87.98 87.00 88.01 70.35 67.40 89.72 72.33 70.14 80.87
SS3D++ (PV-RCNN-based [6]) Sparse 89.32 83.42 78.67 90.23 87.91 87.33 88.33 76.13 70.72 93.99 77.54 72.42 82.89
@@ 5.2 Implementation Details
[…]
${}^{\dagger}$oxel-RCNN [23] Full 76.13 75.66 68.18 67.74 78.20 71.98 69.29 63.59 70.75 69.68 68.25 67.21 70.56
${}^{\dagger}$oxel-RCNN [23] Sparse 47.83 46.32 41.83 40.51 25.88 20.73 21.78 17.44 50.29 48.48 48.38 46.65 38.01
Ours(Voxel-RCNN-based) Sparse 65.99 65.29 57.50 56.90 53.44 48.89 46.41 42.37 60.13 58.93 57.40 56.26 55.79
@@
PV-RCNN++ [56] Sparse 29.96 29.20 25.95 25.29 12.26 10.35 10.20 8.61 43.66 42.54 41.99 40.91 26.74
Ours (PV-RCNN++-based) Sparse 65.73 64.96 57.34 56.67 53.62 48.51 45.33 41.01 59.78 58.36 57.06 55.72 55.34
@@
CenterPoint [63] Sparse 32.98 32.33 28.82 28.24 20.22 16.91 17.30 14.47 28.84 27.90 27.74 26.84 25.22
Ours(CenterPoint-based) Sparse 63.49 62.93 55.30 54.81 48.93 43.89 41.53 37.22 51.87 50.81 50.07 49.04 50.82
@@
5.3 Comparisons with Fully-Supervised Methods
[…]
Our SS3D++ method is detector-agnostic and can be easily used to further improve these off-the-shelf 3D fully-supervised detectors with sparse supervision. Table I shows that our SS3D++ method can substantially boost the detection performance of different fully-supervised detectors for each category across different difficulty levels, and approaches or even surpasses the results of these methods with full supervision. This highlights the significant advantage of the proposed method in eliminating the negative influence of sparse supervision. The visual qualitative analysis of detection results is shown in the right column of Fig. 9. As can be seen, the SS3D++ method effectively overcomes the mistakes of the PV-RCNN detector, and achieves fine detection results. The visualization results of other detectors can be found in Appendix Fig. 1. This further validates the effectiveness of iterations between detector amelioration and confident fully-annotated scene generation.
@@ 5.3 Comparisons with Fully-Supervised Methods
Table II further compares the SS3D++ method with our previous conference method SS3D. The novel SS3D++ method outperforms the previous SS3D method by a margin on three different difficulty levels for all categories. These consistent improvements demonstrate the effectiveness of the mutual benefit of our SS3D++ between 3D detector training and confident fully-annotated scene generation.
@@
[…]
Due to the crowded nature of the Waymo dataset, the distribution of objects within the scenes is usually dense. When we only annotate one instance per scene, it poses significant challenges for detection learning. The large number of missing annotated instances results in seriously incomplete supervision, which causes a detrimental performance. We further attempt to annotate more instances for each scene to improve the performance. Fig. 10 illustrates the varying curve of average detection performance as the number of annotated instances increase per scene. It can be observed that we could achieve about 90% of the performance achieved by fully supervised methods (as depicted by the dashed line in Fig. 10 (a)) with only additionally annotating two objects per scene. Especially, we only require about 15×15$\times$15 × less annotation cost compared with full annotation. We also show the varying curve of detection performance in terms of the vehicle, pedestrian, and cyclist categories in Fig. 10 (b, c, d). It can be seen that the detection performance of different categories can be progressively improved as the number of annotated instances increase per scene. This implies that the SS3D++ method is capable of boosting performance by employing more annotation information.

### v2

@@ 3 Preliminaries
where $E_{i}=\{E_{ij}\},j=1,\cdots,N_{i}$ is a set of detected 3D objects in the $i$-th scene with an input scene $P_{i}\in\mathbb{R}^{N\times 3}$ containing $N$ points, $\phi$ is the weight parameter of $F$, $M$ is the total number of scenes in the dataset, and $E_{ij}$ can be written as
@@ 3.2 Fully-Supervised 3D Object Detection
where classification loss $\mathcal{L}_{cls}$ is used to learn the classifier to predict the category of a 3D object, and $\mathcal{L}_{reg}$ is utilized to learn the location, size and heading angle. If we take the proposal-based two-stage 3D detector Voxel-RCNN [5] as an example, the Eq. (7) degenerates as
@@
where $\mathcal{L}_{cls}^{RPN}$ is the Focal Loss [19] for classification in the region proposal network stage, $\mathcal{L}_{cls}^{Head}$ is the Binary Cross Entropy Loss for confidence prediction in the detection head stage, and Huber Loss is exploited in $\mathcal{L}_{reg}^{RPN}$ and $\mathcal{L}_{reg}^{Head}$ for box regression.
@@ 4 Sparsely-Supervised 3D Object Detection Framework and Learning Algorithm
Though fully-supervised 3D object detection obtains impressive performance, they require large-scale, precisely densely-annotated 3D training data. This data collection process is quite time-consuming and labor-intensive, which prevents the potential applicability of 3D object detection in real-world settings. In this work, we study the following sparsely annotated 3D object detection setting, i.e., we have only access to one 3D bounding box annotation $\tilde{G}_{i}=\{(b_{i1},y_{i1})\}$ for any scene $P_{i}$, and the annotations of remaining object instances are missing. The subsequent experiments verifies that the proposed method is robust in terms of the way of object selection. Such inexact and incomplete supervision greatly saves annotation efforts. In contrast to prior weakly-supervised attempts [68, 29, 63] (as shown in Fig. 1), our novel sparse supervision form introduces more potential practical and commercial benefits, while bringing some new challenges (please refer to Section 1). In this section, we will present the proposed SS3D++ method to address these challenges.
@@ 4.1 Overview SS3D++ Framework and Modeling
$u^{(t+1)}_{ijc}$ $=\left\{\begin{array}[]{ll}1&\mathcal{L}^{c}_{cls}(p_{ij},y_{ij}^{(t)};\phi^{(t)})<\lambda^{u}_{ic}\\ 0&\mathcal{L}^{c}_{cls}(p_{ij},y_{ij}^{(t)};\phi^{(t)})\geq\lambda^{u}_{ic}\end{array},\right$. (16)
$v^{(t+1)}_{ijc}$ $=\left\{\begin{array}[]{ll}1&\mathcal{L}^{c}_{reg}(\mathcal{A}(r_{ij}),\tilde{r}_{ij};\phi^{(t)})<\lambda^{v}_{ic}\\ 0&\mathcal{L}^{c}_{reg}(\mathcal{A}(r_{ij}),\tilde{r}_{ij};\phi^{(t)})\geq\lambda^{v}_{ic}\end{array},\right$. (19)
$k^{(t+1)}_{ijc}$ $=\left\{\begin{array}[]{ll}1&d^{c}(r_{ij};\phi^{(t)})<\lambda^{k}_{ic}\\ 0&d^{c}(r_{ij};\phi^{(t)})\geq\lambda^{k}_{ic}\end{array},\right$. (22)
@@
Algorithm 1 Our SS3D++ Method.
@@
${\phi}_{s}^{*}=\mathop{\arg\min}_{{\phi}}(\mathcal{L}_{s}+\mathcal{L}_{p}).$ (24)
$\phi_{m}\leftarrow\alpha\phi_{m}+(1-\alpha)\phi_{s},$ (26)
where $\alpha$ is the EMA coefficient. Note that the training data are the union set of initial sparse annotated instances and selected instances ($u_{ij}=v_{ij}=k_{ij}=1$) with the pseudo annotation $\tilde{U}_{i}$. To increase the learning efficiency of the detector, we further generate confident fully-annotated scenes based on the mined reliable background and confident missing-annotated instances. The details can refer to Sec. 4.4. Thus, this step can be solved by the existing standard off-the-shelf 3D object detector, described as PointRCNN [44], Part-A2[45], PV-RCNN[42], Voxel-RCNN[5], etc.
@@ 4.2.2 Adaptive Sample Selection Curriculum Setups
Figure 7: Illustration of the proposed reliable background mining module. To start, we feed the original point cloud to the detector without NMS to produce near-duplicate predictions, and leverage the instances stored in the instance bank to filter out unreliable 3D object points. This will lead to broken scene with reliable background, which is further processed through the GT sampling data augmentation in Sec. 4.4.
Algorithm 2 Confident Missing-Annotated Instance Mining.
@@
3:  Obtain ${E}_{i}$ and $\hat{{E}}_{i}$ by inputting $P_{i}$ and $\hat{P}_{i}$ to the detector $F_{m}(\cdot;\phi^{(t)})$ based on Eq. (5);
4:  Set $\lambda^{u}_{ic},\lambda^{v}_{ic},\lambda^{v}_{ic}$ by adaptive sample selection curriculum setups in Eq. (4.2.2);
5:  Perform joint multi-criteria sample selection process via Eq. (16), Eq. (19) and Eq. (22);
6:  Generate high quality of pseudo annotations $\tilde{U}_{i}$ for selected instances via Eq. (23);
@@ 4.3 Reliable Background Mining
As illustrated in Fig. 7, we construct an instance bank which initially just stores the sparsely annotated instances and then progressively stores mined highly confident pseudo instances (please see in Section 4.2).
Specifically, we denote the predicted 3D bounding boxes as $R_{i}$, and the bounding boxes of the scene $P_{i}$ in instance bank $\mathcal{B}$ as $\mathcal{B}_{i}$, which contains the sparsely annotated instance and previously mined pseudo instances. And then we delete all points of the predicted bounding boxes $R_{i}$ in the scene $P_{i}$ to get the confident scene $\tilde{P}_{i}$. Next, the bounding boxes and corresponding points stored in the instance bank $\mathcal{B}_{i}$ of the $i$-th scene will fill into the scene $P_{i}$. The instance bank helps to avoid wrongly discarding reliable 3D objects whilst removing the ambiguous points for each training scene.
@@
Algorithm 3 Reliable Background Mining.
@@ 5.2 Implementation Details
Method Data 3D Detection (Car) BEV Detection (Car) 3D Detection (Cyc.) BEV Detection (Cyc.) Avg.
Easy Mod Hard Easy Mod Hard Easy Mod Hard Easy Mod Hard
SS3D (PointRCNN-based) Sparse (20%) 87.18 77.10 76.13 89.74 87.41 85.71 86.62 73.22 66.92 87.21 74.27 71.54 80.25
SS3D++ (PointRCNN-based) Sparse (20%) 88.63 77.90 77.13 90.26 87.65 86.10 86.73 74.81 70.28 87.76 74.64 71.81 81.14
Improvements - +1.45 +0.80 +1.00 +0.52 +0.24 +0.39 +0.11 +1.59 +3.36 +0.55 +0.37 +0.27 +0.89
SS3D (PV-RCNN-based) Sparse (20%) 89.49 79.30 78.28 90.45 87.98 87.00 88.01 70.35 67.40 89.72 72.33 70.14 80.87
SS3D++ (PV-RCNN-based) Sparse (20%) 89.32 83.42 78.67 90.23 87.91 87.33 88.33 76.13 70.72 93.99 77.54 72.42 82.89
Improvements - -0.17 +4.12 +0.39 -0.22 -0.07 +0.33 +0.32 +5.78 +3.32 +4.27 +5.21 +2.28 +2.02
5.3 Comparisons with Fully-Supervised Methods
Comparison with fully-supervised methods on KITTI. To evaluate our SS3D++ framework on the highly-competitive KITTI dataset, we compare the proposed method with four SOTA fully-supervised methods: PointRCNN [44], Part-A2 [45], PV-RCNN [42], Voxel-RCNN [5], with fully-annotated train split and the extremely sparse train split, respectively, where these detectors trained on the extremely sparse split are used as the initial detectors of our method. The results of different methods are shown in Tab. I. From this table, it can be seen that the four fully-supervised detectors trained on the sparsely annotated dataset produce a significant performance degradation due to the extremely inexact and incomplete supervision. For example, PointRCNN has an average performance decrease of approximately 25% points on “car” 3D detection, 21% points on “pedestrian” 3D detection, and 25% points on “cyclist” 3D detection. Even for the competitive PV-RCNN detector, it exhibits an average decline of about 19% points on “car” 3D detection, 6% points on “pedestrian” 3D detection, and 8% points on “cyclist” 3D detection. In the left column of Fig. 9, we visualize the detection results of PV-RCNN on the KITTI dataset. As shown, there exist some typical mistakes for the PV-RCNN detector, e.g., the background wrongly detected as an object (indicated by blue arrows), missing object detection (pink arrows), inaccurate localization (yellow arrows), etc.
@@ 5.3 Comparisons with Fully-Supervised Methods
PartA2[45] Full 74.66 74.12 65.82 65.32 71.71 62.24 62.46 54.06 66.53 65.18 64.05 62.75 65.74
PartA2[45] Sparse (2.2%) 43.95 42.40 38.35 37.00 24.27 19.51 20.37 16.36 49.92 48.46 48.02 46.61 36.27
Ours (PartA2-based) Sparse (2.2%) 66.27 65.35 58.03 57.02 52.54 45.68 44.37 38.46 60.71 58.80 58.12 56.67 55.17
Voxel-RCNN [5] Full 76.13 75.66 68.18 67.74 78.20 71.98 69.29 63.59 70.75 69.68 68.25 67.21 70.56
Voxel-RCNN [5] Sparse (2.2%) 47.83 46.32 41.83 40.51 25.88 20.73 21.78 17.44 50.29 48.48 48.38 46.65 38.01
Ours(Voxel-RCNN-based) Sparse (2.2%) 65.99 65.29 57.50 56.90 53.44 48.89 46.41 42.37 60.13 58.93 57.40 56.26 55.79
@@
PV-RCNN++ [43] Sparse (2.2%) 29.96 29.20 25.95 25.29 12.26 10.35 10.20 8.61 43.66 42.54 41.99 40.91 26.74
Ours (PV-RCNN++-based) Sparse (2.2%) 65.73 64.96 57.34 56.67 53.62 48.51 45.33 41.01 59.78 58.36 57.06 55.72 55.34
@@
CenterPoint [78] Sparse (2.2%) 32.98 32.33 28.82 28.24 20.22 16.91 17.30 14.47 28.84 27.90 27.74 26.84 25.22
Ours(CenterPoint-based) Sparse (2.2%) 63.49 62.93 55.30 54.81 48.93 43.89 41.53 37.22 51.87 50.81 50.07 49.04 50.82
Our SS3D++ method is detector-agnostic and can be easily used to further improve these off-the-shelf 3D fully-supervised detectors with sparse supervision. Tab. I shows that our SS3D++ method can substantially boost the detection performance of different fully-supervised detectors for each category across different difficulty levels, and approaches or even surpasses the results of these methods with full supervision. This highlights the significant advantage of the proposed method in eliminating the negative influence of sparse supervision. The visual qualitative analysis of detection results is shown in the right column of Fig. 9. As can be seen, the SS3D++ method effectively overcomes the mistakes of the PV-RCNN detector, and achieves fine detection results. The visualization results of other detectors can be found in Appendix Fig. 18. This further validates the effectiveness of iterations between detector amelioration and confident fully-annotated scene generation.
@@
Tab. II further compares the SS3D++ method with our previous conference method SS3D. The novel SS3D++ method outperforms the previous SS3D method by a margin on three different difficulty levels for all categories. These consistent improvements demonstrate the effectiveness of the mutual benefit of our SS3D++ between 3D detector training and confident fully-annotated scene generation.
TABLE IV: Comparison with other sparsely-supervised methods trained with extremely sparse annotated KITTI dataset. The 3D object detection benchmark is evaluated by mean average precision with R40, under IoU thresholds 0.7.
Method Venue Data Car - 3D Detection Car - BEV Detection Avg.
Easy Mod Hard Easy Mod Hard
Voxel-RCNN [5] AAAI’21 Sparse (2%) 72.5 54.9 44.8 83.6 71.4 57.7 64.15
Coln [72] ICCV’23 Sparse (2%) 84.5 68.4 58.0 92.3 81.0 70.2 75.73
Coln++ [72] ICCV’23 Sparse (2%) 92.0 79.5 71.5 96.1 88.8 82.5 85.06
SP3D [85] CVPR’25 Sparse (2%) 91.3 74.8 63.8 95.4 85.2 74.5 80.83
SS3D++ (Ours) - Sparse (2%) 93.2 80.7 75.9 96.6 90.0 84.6 86.83
TABLE V: Comparison with Coln trained with extremely sparse annotated Waymo Open Dataset.
Veh. (LEVEL 1) Veh. (LEVEL 2) Ped. (LEVEL 1) Ped. (LEVEL 2) Cyc. (LEVEL 1) Cyc. (LEVEL 2)
Methods Data mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH Avg.
CenterPoint [78] Fully 71.33 70.76 63.16 62.65 72.09 65.49 64.27 58.23 68.68 67.39 66.11 64.87 66.25
CenterPoint (From Coln++) Sparse (2.2%) 32.15 31.55 27.97 27.45 25.66 21.65 22.00 18.56 59.25 57.84 57.22 55.86 36.43
Coln++ [72] Sparse (2.2%) 48.25 47.60 41.82 41.25 28.25 24.28 23.79 20.45 63.99 62.60 61.71 60.37 43.69
Improvements - +16.10 +16.05 +13.85 +13.80 +2.59 +2.63 +1.79 +1.89 +4.74 +4.76 +4.49 +4.51 +7.26
CenterPoint [78] (Our produced) Sparse (2.2%) 32.98 32.33 28.82 28.24 20.22 16.91 17.30 14.47 28.84 27.90 27.74 26.84 25.22
SS3D++ (Ours) Sparse (2.2%) 63.49 62.93 55.30 54.81 48.93 43.89 41.53 37.22 51.87 50.81 50.07 49.04 50.82
Improvements - +30.51 +30.60 +26.48 +26.57 +28.71 +26.98 +24.23 +22.75 +23.03 +22.91 +22.33 +22.20 +25.61
TABLE VI: Performance comparison with state-of-the-art semi-supervised methods on the Waymo Open Dataset with 202 validation sequences for the 3D detection. Bold text and underlined text represent the best performance and the second-best performance, respectively. The “Improvement” indicates the difference between the best and the second best.
1% Annotation Data Venue Veh. (LEVEL 1) Veh. (LEVEL 2) Ped. (LEVEL 1) Ped. (LEVEL 2) Cyc. (LEVEL 1) Cyc. (LEVEL 2) Avg.
mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH
PV-RCNN [42] CVPR’20 48.5 46.2 45.5 43.3 30.1 15.7 27.3 15.9 4.5 3.0 4.3 2.9 23.93
DetMatch [32] ECCV’22 52.2 51.1 48.1 47.2 39.5 18.9 35.8 17.1 - - - - 38.74
HSSDA [22] CVPR’23 56.4 53.8 49.7 47.3 40.1 20.9 33.5 17.5 29.1 20.9 27.9 20.0 34.76
MT [17] ICCV’25 59.8 57.7 52.0 50.2 44.0 22.7 37.5 19.3 35.2 19.5 33.9 18.7 37.54
A-Teacher [62] CVPR’24 56.5 54.5 49.2 47.5 48.1 27.3 40.8 23.1 35.1 27.1 33.7 26.1 39.08
PPTM [70] AAAI’24 61.5 59.8 53.7 52.2 43.1 22.3 36.3 18.8 35.7 17.9 35.7 34.3 39.28
SS3D++ (Ours) - 62.1 61.3 54.2 53.5 50.9 45.4 43.0 38.4 52.8 51.5 50.8 49.5 51.12
Improvements - +0.6 +1.5 +0.5 +1.3 +2.8 +18.1 +2.2 +15.3 +17.1 +24.4 +15.1 +15.2 +9.51
Comparison with fully-supervised methods on Waymo. The Waymo dataset is larger than the KITTI dataset and the scene is very crowded, with an average of 44 objects in each scene. Thus it is relatively difficult and expensive to provide full precise annotations. The proposed framework of using extremely sparse annotation can effectively reduce the workload to a greater extent (about 2.2% points). However, such sparse annotation significantly deteriorates the existing fully-supervised detectors. As shown in Tab. III, PartA2 and Voxel-RCNN have an average performance decrease of beyond 29% and 32% points, respectively. To further verify the generalizability of our SS3D++, we also replace the base detector with two latest networks that are different from those in the KITTI experiment, i.e., PV-RCNN++[43] and CenterPoint [78]. While they encounter an average performance decrease of beyond 44% and 41% points, respectively.
Nonetheless, our SS3D++ method can still achieve about 80% performance compared with the detectors trained with full supervision and improves the average detection performance of PartA2, Voxel-RCNN, PV-RCNN++, and CenterPoint by 18.90%, 17.76%, 28.60%, 25.60% points, respectively. This demonstrates that SS3D++ method can achieve a consistent performance improvement across different datasets/detectors. Observing that we only have access to only 2.2% precise annotations, this implies that the SS3D++ method is hopeful to be readily applied to real-world 3D object detection that requires low-cost annotations. The qualitative comparison results on the Waymo Open dataset can be found in Appendix Fig. 18.
Due to the crowded nature of the Waymo dataset, the distribution of objects within the scenes is usually dense. When we only annotate one instance per scene, it poses significant challenges for detection learning. The large number of missing annotated instances results in seriously incomplete supervision, which causes a detrimental performance. We further attempt to annotate more instances for each scene to improve the performance. Fig. 10 illustrates the varying curve of average detection performance as the number of annotated instances increase per scene. It can be observed that we could achieve about 90% of the performance achieved by fully supervised methods (as depicted by the dashed line in Fig. 10 (a)) with only additionally annotating two objects per scene. Especially, we only require about $15\times$ less annotation cost compared with full annotation. We also show the varying curve of detection performance in terms of the vehicle, pedestrian, and cyclist categories in Fig. 10 (b, c, d). It can be seen that the detection performance of different categories can be progressively improved as the number of annotated instances increase per scene. This implies that the SS3D++ method is capable of boosting performance by employing more annotation information.
@@
5.4 Comparisons with SOTA Methods
Comparison with the sparsely-supervised method. We have compared our method with recently SOTA methods [72, 85, 88] under the sparsely-supervised setting. Tab. IV presents the performance comparison on the KITTI dataset. To ensure a fair comparison, we used Voxel-RCNN as the backbone detector with 2% labeled data and reported the Recall@40 performance at an IoU threshold of 0.7. It is evident that our method outperforms Coln++ across different difficulty levels. Moreover, we achieved the highest average performance across all methods. Additionally, Tab. V provides the performance comparison on the Waymo dataset [57] under the same sparse annotation settings. Similarly, it can be observed that our SS3D++ significantly outperforms Coln++ in this setting as well, highlighting its robustness and generalization capability across datasets.
Comparison with the semi-supervised method. First, we compare our results on the large-scale Waymo dataset [57] with current state-of-the-art semi-supervised methods. For fair comparison, we keep all methods trained with the same number of annotated objects and our SS3D++ adopts PV-RCNN as the backbone detector. Remarkably, with just 1% of the annotation effort, we are able to achieve competitive results, as illustrated in Tab. VI. As can be seen from the last row, our method outperforms the semi-supervised methods with best performance across all categories, with an average improvement of 9.5 points.